ss19.2.9

In an ideal thermal engine, 300 J is generated with each kilojoule of energy absorbed by the warmer reservoir. The temperature of the colder reservoir is 280 K. What is the temperature of the warmer reservoir? (Carnot Process)
Answer #1

Solution -

Efficiency = work / heat input = 300 / 1000 = 0.3

Efficiency = 1 - T2 / T1   

0.3 =   1 - T2 / T1   

T2 / T1 = 0.7

280  / T1 = 0.7

T1 =   280 / 0.7 = 400 K

Please upvote , if you are satisfied with solution.